# A QUBO Formulation for the Generalized LinkedIn Queens Game

### v1

@@ Abstract
In this paper, we present a QUBO formulation designed to solve a series of generalisations of the LinkedIn queens game, a version of the N-queens problem. We adapt this formulation for several particular cases of the problem by trying to optimise the number of variables and interactions, improving the possibility of applying it on quantum hardware by means of Quantum Annealing or the Quantum Approximated Optimization Algorithm (QAOA). We also present two new types of problems, the Coloured Chess Piece Problem and the Max Chess Pieces Problem, with their corresponding QUBO formulations.
@@ 1 Introduction
LinkedIn recently published its Queens game (which we will call LQueens), based on N-queens problem, but with some changes in the restrictions to be carried out and the initial conditions. In this new version, the restriction on diagonals is eased, but a restriction on regions is imposed.
In recent years, quantum computing has gained great popularity due to its different computational capabilities compared to classical computing, making it possible to tackle certain problems more efficiently. Some of these problems are combinatorial optimisation problems, in which quantum annealing using quantum annealers [2, 3] such as those of Dwave and the Quantum Approximated Optimization Algorithm (QAOA) [4] using digital quantum computers stand out. In both cases, the formulation of the problem comes through a quadratic unconstrained optimisation problem (QUBO) [5], which contains the interactions to be optimised between the different variables, represented by logical qubits, minimising the energy of the system to find the lowest cost combination. These methods can be adapted to cases with constraints by adding terms that increase the energy substantially if the constraints are not respected.
Recently, the N-queens problem has been solved with a quantum algorithm in a quantum simulator [6], but not generalized and formulated in an optimal formulation for the currently most standardized algorithms. On the other hand, there are also QUBO formulations for other games such as Sudoku [7].
Motivated by this new version of the game and the quantum computing algorithms, we present a series of generalisations of the N-queens problem and a QUBO formulation associated with each of them so that they can be solved by quantum means, such as quantum annealing or the Quantum Approximated Optimization Algorithm (QAOA). With these generalizations, we can define two new types of problems, which take advantage of the definitions and terms we created for the N-queens generalizations.
@@ 3 QUBO resolution
For the LQueens the formulation is similar. The first adjustment we have to make is to consider what are the fixed variables due to the initial condition. These are removed directly from the optimisation problem, but we keep the notation of the boxes for clarity. That is, if we have the $(3,3)$ box, the variable referring to the $(4,3)$ box will still be $x_{4,3}$, since we know that there really is no $x_{3,3}$. So we will have a number of variables equal to $N^{2}$ if there are no queens at the beginning, and less if we add queens in the initial condition.
@@ 4 General cases
4.1 Number of queens per row/column
@@
4.2 Distant diagonals constraint
@@
4.3 Number of queens per region
@@
4.4 Two possible numbers of queens per region
@@
As can be seen, the diagonal constraints are only particular cases of this constraint in which there is one region for each cell, which would be its diagonals. The value $v_{k}$ would be 0 for all the regions, since the number of queens must be 0 or 1. We can rewrite the eq (4.3), with $N_{R}$ regions $R_{k}$ with a single possible number of queens and $N_{V}$ regions $V_{k}$ with two consecutive possible numbers of queens, as
@@
5 Definition of generalised N-queens
@@
5.1 N-queens with regions
@@
5.2 N-queens with soft regions
@@
6 Chess Piece Problems
@@
Given a board of arbitrary shape, of maximum height $N$ and maximum width $M$, on which there are placed a number $N_{p}$ of initial pieces in positions given by the set $A_{I}$, and divided into $N_{R}$ regions, which cannot overlap each other, such that in each region there must be only one piece, the objective will be to place all the possible pieces on the board such that no piece threatens another. Each cell $(i,j)$ can be empty or have the piece of type $P_{ij}$, with possible moves $M_{P_{ij}}$, and can reach cells $C\left(M_{P_{ij}})_{ij}\right)$. A solution of this problem is represented in Fig. 4 a.

### v2

# A QUBO Formulation for the Generalized LinkedIn Queens and Takuzu/Tango Game

@@ Abstract
In this paper, we present a QUBO formulation designed to solve a series of generalisations of the LinkedIn queens game, a version of the N-queens problem, for the Takuzu game (or Binairo), for the most recent LinkedIn game, Tango, and for its generalizations. We adapt this formulation for several particular cases of the problem, as Tents & Trees, by trying to optimise the number of variables and interactions, improving the possibility of applying it on quantum hardware by means of Quantum Annealing or the Quantum Approximated Optimization Algorithm (QAOA). We also present two new types of problems, the Coloured Chess Piece Problem and the Max Chess Pieces Problem, with their corresponding QUBO formulations.
@@ 1 Introduction
Combinatorial games have always been an inspiration for developing and perfecting new mathematical and computational problem-solving techniques, as well as for benchmarking. Combinatorial problems are known for their practical application in academic and industrial contexts, but also for their usefulness in creating entertaining games. Some of these are the Akari, the KenKen, the Takuzu or the N-queens, which led to different combinatorial games, among which is the LinkedIn Queens.
@@
LinkedIn recently published its Queens game (which we will call LQueens), based on N-queens problem, but with some changes in the restrictions to be carried out and the initial conditions. In this new version, the restriction on diagonals is eased, but a restriction on regions is imposed. Following the success that this game has had among the community, LinkedIn announced its new game called Tango, a version of the Takuzu game (also named Binairo or 0h h1) for a $6\times 6$ board, which consists of placing a set of suns and moons on a board following a series of constraints. This game can be visualized as a single-player version of tic-tac-toe, and has some applications such as the modeling of quantum agents [2]. This problem has been solved previously using genetic algorithms [3].
In recent years, quantum computing has gained great popularity due to its different computational capabilities compared to classical computing, making it possible to tackle certain problems more efficiently. Some of these problems are combinatorial optimisation problems, in which quantum annealing using quantum annealers [4, 5] such as those of Dwave and the Quantum Approximated Optimization Algorithm (QAOA) [6] using digital quantum computers stand out. In both cases, the formulation of the problem comes through a quadratic unconstrained optimisation problem (QUBO) [7], which contains the interactions to be optimised between the different variables, represented by logical qubits, minimising the energy of the system to find the lowest cost combination. These methods can be adapted to cases with constraints by adding terms that increase the energy substantially if the constraints are not respected. An example can be the QUBO formulation of the Sudoku game [8].
Recently, the N-queens problem has been solved with a quantum algorithm in a quantum simulator [9], with two quantum algorithms based on the W-states and backtracking [10], and has also been solved with several QUBO formulations [11, 12]. However, these solutions are tailored and restricted to the N-queens, and not to all generalizations we may have of the problem.
Motivated by this new version of the game and the quantum computing algorithms, we present a series of generalisations of the N-queens problem and a QUBO formulation associated with each of them so that they can be solved by quantum means, such as quantum annealing or the Quantum Approximated Optimization Algorithm (QAOA). With these generalizations, we can define two new types of problems, which take advantage of the definitions and terms we created for the N-queens generalizations. Due to the popularity of Takuzu as a combinatorial problem and its connection to quantum computing, it is also interesting to propose a QUBO formulation for Takuzu, Tango and generalizations of it that we want to formulate.
@@ 2.2 QUBO resolution
For the LQueens the formulation is similar. The first adjustment we have to make is to consider what are the fixed variables due to the initial condition. These are removed directly from the optimisation problem, but we keep the notation of the celles for clarity. That is, if we have the $(3,3)$ cell, the variable referring to the $(4,3)$ cell will still be $x_{4,3}$, since we know that there really is no $x_{3,3}$. So we will have a number of variables equal to $N^{2}$ if there are no queens at the beginning, and less if we add queens in the initial condition.
@@ 2.3 General cases
2.3.1 Number of queens per row/column
@@
2.3.2 Distant diagonals constraint
@@
2.3.3 Number of queens per region
@@
2.3.4 Two possible numbers of queens per region
@@
As can be seen, the diagonal constraints are only particular cases of this constraint in which there is one region for each cell, which would be its diagonals. The value $v_{k}$ would be 0 for all the regions, since the number of queens must be 0 or 1. We can rewrite the eq (2.3.3), with $N_{R}$ regions $R_{k}$ with a single possible number of queens and $N_{V}$ regions $V_{k}$ with two consecutive possible numbers of queens, as
@@
2.3.5 Toroidal board
An interesting generalization is to consider a toroidal board, so that the pieces that pass through the upper part of the board exit through the lower part, those that pass through the right part exit through the left part and vice versa. To consider this problem, we must first impose that the side of the board has a minimum prime factor of 5, otherwise it will have no solution. We can also approach the case of a cylindrical board, in which this cyclic behavior will only occur when crossing the left and right sides.
Figure 4: Toroidal board with a queen, where we mark in red what would be the squares that this queen could attack from only one of its 4 diagonals.
To take into account this generalization, the terms of rows and columns will be exactly the same, as well as the term of the regions. The only term that will change will be that of the diagonals, since this, according to the length and shapes of the diagonals, will have to consider the cells reached after curving the boundary as many times as appropriate. Therefore, we will only have to change the definition of the $D^{\prime}_{ij}$ to include all the necessary cells.
2.4 Definition of generalised N-queens
@@
2.4.1 N-queens with regions
@@
2.4.2 N-queens with soft regions
@@
2.5 Tents & Trees
An interesting game related to these generalizations is Tents & Trees. This game is formulated as follows: Given a board of dimensions $N\times N$ cells, with $N_{t}$ trees placed at positions $A_{t}$, the objective is to place a series of tents such that the following constraints are satisfied:
  • Each tree will have an associated tent, in a cell vertically or horizontally adjacent to it.
  • No two tents can be adjacent to each other, not even diagonally.
  • Each row $i$ and column $j$ has to have a number $N_{r,i}$ and $N_{c,j}$ of associated tents.
We can translate this problem to our formalism with the following considerations:
  • Each tree cell is no longer considered as an optimizing variable and is fixed to 0.
  • The $k$-th tree creates a region $R_{k}$, given by the 4 cells it has adjacent to it, such that in this region there must be at least 1 tent. In the cells not included by any of the regions, we fix the variables to 0.
  • The tents act as queens, but can only attack the 8 adjacent cells.
  • Each row $i$ and column $j$ has to have a number $N_{r,i}$ and $N_{c,j}$ of associated tents.
Our variables will be a set of binary $x_{ij}$ indicating that in row cell $i$ and column $j$ there is no tent if equal to 0 and there is if equal to 1.
The terms imposing the number of tents per row and column are exactly as we posed in the general case in (10) and (11). For the non-contact constraint between tents, we are going to use the term of (12), but making $D^{\prime}_{i,j}$, in addition to taking into account the adjacent diagonals, take into account the vertical and horizontal adjacencies. To reduce the number of terms, we will not take into account the cells above the tent, because if a tent has another one above it, the one above it will see it below it.
For the regions given by the trees, we will have to use a term of type (15) with $v_{k}=1$. This is because if there were 3 or 4 tents surrounding the tree, there would necessarily be 2 of them in diagonal contact. Therefore, there can only be 1 or 2 tents in that region.
Thus, the QUBO of the problem is
$Q=$ $Q_{r}+Q_{c}+Q_{d}+Q_{g}=$
$=$ $\sum_{i=1}^{N}\left(N_{r,i}-\sum_{j=1}^{N}x_{ij}\right)^{2}+\sum_{j=1}^{N}\left(N_{c,j}-\sum_{i=1}^{N}x_{ij}\right)^{2}+$
$+$ $\sum_{i,j}^{N,N}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl}+$
$+$ $\sum_{k=1}^{N_{t}}\left(\frac{3}{2}-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2}.$ (22)
As in the previous cases, the minimum energy will be of $\frac{N_{t}}{4}$.
2.6 Chess Piece Problems
@@
Given a board of arbitrary shape, of maximum height $N$ and maximum width $M$, on which there are placed a number $N_{p}$ of initial pieces in positions given by the set $A_{I}$, and divided into $N_{R}$ regions, which cannot overlap each other, such that in each region there must be only one piece, the objective will be to place all the possible pieces on the board such that no piece threatens another. Each cell $(i,j)$ can be empty or have the piece of type $P_{ij}$, with possible moves $M_{P_{ij}}$, and can reach cells $C\left(M_{P_{ij}}\right)_{ij}$. A solution of this problem is represented in Fig. 5 a.
@@
3 Takuzu and LinkedIn’s Tango Problems
3.1 Problem definition
The problem to be solved in Takuzu is the following:
Given a square board of dimension $N\times M$ cells, where each cell can contain only one 0 or 1, initially with $N_{I0}$ zeros placed at positions $A_{I0}$ and $N_{I1}$ ones placed at positions $A_{I1}$, we have to determine for each cell whether there should be a zero or a one. The constraints are as follows:
  • In each row and column there must be as many zeros as ones.
  • There cannot be more than two zeros or ones in a row vertically or horizontally.
  • No two rows and no two columns can have the same combination of zeros and ones.
A solution to this problem can be seen in Fig. 6 a.
The Tango problem is very similar, but adding additional constraints and removing the restriction of repeating rows or columns. It is expressed as follows:
Given a square board of dimension $6\times 6$ cells, where each cell can contain only one sun icon or one moon icon, initially with $N_{IS}$ suns placed at positions $A_{IS}$ and $N_{IM}$ moons placed at positions $A_{IM}$, we have to determine for each cell whether there should be a sun or a moon. The constraints are as follows:
  • In each row and column there must be as many moons as suns.
  • There cannot be more than two suns or moons in a row vertically or horizontally.
  • There are cells between which there is an ’=’ symbol, which indicates that in both cells there must be the same icon.
  • There are cells between which there is an ’x’ symbol, which indicates that there must be a different icon in both cells.
A solution to this problem can be seen in Fig. 6 b.
Figure 6: Solution for a problem a) Takuzu/0h h1 $6\times 6$, b) Tango
In this case, we can see that if we replace the suns by zeros and the moons by ones, and allow the board to be $N\times M$, we will have a generalization of Takuzu. We call this problem the Takuzu/Tango Problem (TTP).
To formulate this problem, the state of a board is given by a set of variables $x_{ij}$, which indicate the value in the cell of row $i$ and column $j$. Therefore, the constraints of the problem can be expressed as:
  • Row regularity: there must be as many zeros as ones in each row (Fig. 7 c):
    $\sum_{j=1}^{M}x_{ij}=\frac{M}{2}\quad\forall i\in[1,N].$ (28)
  • Column regularity: there must be as many zeros as ones in each column (Fig. 7 c):
    $\sum_{i=1}^{N}x_{ij}=\frac{N}{2}\quad\forall j\in[1,M].$ (29)
  • Vertical repetition condition: there cannot be more than two zeros or ones in a row vertically. Therefore, for each group of 3 vertically consecutive cells, there can only be 1 or 2 ones, which implies that their associated variables can only add up to 1 or 2 (Fig. 7 a):
    $\sum_{k=i}^{i+2}x_{kj}\in\{1,2\}\quad\forall i\in[1,N-2],j\in[1,M].$ (30)
  • Horizontal repetition condition: there cannot be more than two zeros or ones in a row horizontally. Therefore, for each group of 3 horizontally consecutive cells, there can only be 1 or 2 ones, which implies that their associated variables can only add up to 1 or 2 (Fig. 7 b):
    $\sum_{k=j}^{j+2}$ $x_{ik}\in\{1,2\}$ (31)
    $\forall i$ $\in[1,N],j\in[1,M-2].$
  • Equal sign condition: If there is an ’=’ symbol between two cells, both cells must have the same value. These symbols can be between two cells horizontally or between two cells vertically (Fig. 7 d). There are $N_{He}$ horizontal symbols and $N_{Ve}$ vertical symbols. The $k$-th horizontal ’=’ is between positions $(I^{He}_{k},J^{He}_{k})$ and $(I^{He}_{k},J^{He}_{k}+1)$, while the $k$-th vertical ’=’ is between positions $(I^{Ve}_{k},J^{Ve}_{k})$ and $(I^{Ve}_{k}+1,J^{Ve}_{k})$. Therefore, for the horizontals, the following is satisfied
    $x_{I^{He}_{k},J^{He}_{k}}=x_{I^{He}_{k},J^{He}_{k}+1}\ \forall k\in[1,N_{He}],$ (32)
    and for the verticals
    $x_{I^{Ve}_{k},J^{Ve}_{k}}=x_{I^{Ve}_{k}+1,J^{Ve}_{k}}\quad\forall k\in[1,N_{Ve}].$ (33)
  • Cross sign condition: If there is an ’x’ symbol between two cells, both cells must have different value. These symbols can be between two cells horizontally or between two cells vertically (Fig. 7 e). There are $N_{Hc}$ horizontal symbols and $N_{Vc}$ vertical symbols. The $k$-th horizontal ’x’ is between positions $(I^{Hc}_{k},J^{Hc}_{k})$ and $(I^{Hc}_{k},J^{Hc}_{k}+1)$, while the $k$-th vertical ’x’ is between positions $(I^{Vc}_{k},J^{Vc}_{k})$ and $(I^{Vc}_{k}+1,J^{Vc}_{k})$. Therefore, for the horizontals, the following is satisfied
    $x_{I^{Hc}_{k},J^{Hc}_{k}}\neq x_{I^{Hc}_{k},J^{Hc}_{k}+1}\ \forall k\in[1,N_{Hc}],$ (34)
    and for the verticals
    $x_{I^{Vc}_{k},J^{Vc}_{k}}\neq x_{I^{Vc}_{k}+1,J^{Vc}_{k}}\quad\forall k\in[1,N_{Vc}].$ (35)
  • Global non-repetition constraint: no two rows can have the same combination of zeros and ones. As the positions of the ones condition those of the zeros, we will only have to consider those of the ones. If two cells in the same row are at one, the product of their values will be one. In any other case it will be zero. Therefore, all the elements of two rows will coincide if the sum of their products is equal to the number of ones in each row.
    $\sum_{j=1}^{M}x_{aj}x_{bj}<\frac{M}{2}\ \forall a,b\in[1,N]\ |\ a\neq b.$ (36)
    Neither can there be two columns with the same combination of zeros and ones.
    $\sum_{i=1}^{N}x_{ia}x_{ib}<\frac{N}{2}\ \forall a,b\in[1,M]\ |\ a\neq b.$ (37)
Figure 7: Boards with restrictions. The blue cells have the values previously set, while the red cells have their values determined by the constraints. a) Vertical repetition, b) Horizontal repetition, c) Regularity, d) Symbol ’=’, e) Symbol ’x’.
3.2 QUBO Formulation
The QUBO formulation of this problem is composed of different terms that are responsible for imposing each of the constraints. Before starting to formulate each term, we will discuss how to optimize the number of variables to be determined by fixing trivial values for some variables given an initialization. This way, some variables will have a fixed value in the QUBO terms, but they will not be introduced as variables to be optimized in algorithms such as QAOA or Quantum Annealers, but will be part of the coefficients of the cost matrix. This make it possible to reduce the resources, e.g. qubits, to be used in quantum devices.
If we have a set of zeros and ones placed in their positions $A_{I0}$ and $A_{I1}$, the variables associated with those cells do not need to be determined. This will reduce the number of variables to optimize to $NM-N_{I0}-N_{I1}$.
In case any of these already fixed variables is adjacent to an ’=’ symbol, we will trivially fix the adjacent variable to its same value. If the adjacent symbol is an ’x’, we will fix the adjacent variable to the opposite value. This allow us to remove those variables from the optimization in the same way.
In case a row or column has already fixed as many zeros or ones as it should have by the regularity condition, the rest of the variables are set to the opposite value, as this is the only way to satisfy this constraint. With this we can eliminate even more variables trivially.
In case there are 2 adjacent variables fixed with the same value, we will automatically know that the ones adjacent to both of them that follow in their line have to have the opposite value, which allows us to fix some more variables.
Once the number of variables to be determined has been optimized, we proceed with the creation of the QUBO terms. The first term we need to create is the horizontal repetition term, which is
$Q_{HR}=\sum_{i,j=1}^{N,M-2}\left(\frac{3}{2}-\sum_{k=j}^{j+2}x_{ik}\right)^{2}$ (38)
where we have that the minimum is centered between 1 and 2 using a fractional QUBO term [13], so that in each horizontal group of 3 variables the minimum value of the energy ($\frac{1}{4}$) is given when 1 or 2 variables have the value 1. For the vertical condition, we have an analogous term
$Q_{VR}=\sum_{i,j=1}^{N-2,M}\left(\frac{3}{2}-\sum_{k=i}^{i+2}x_{kj}\right)^{2}.$ (39)
For the regularization conditions, we only need a term that imposes that the sum of the variables per row or column results in a specific value. These terms are for each case
$Q_{Hr}=\sum_{i=1}^{N}\left(\frac{M}{2}-\sum_{j=1}^{M}x_{ij}\right)^{2},$ (40)
$Q_{Vr}=\sum_{j=1}^{M}\left(\frac{N}{2}-\sum_{i=1}^{N}x_{ij}\right)^{2}.$ (41)
Finally, we will have the 4 terms necessary to impose the constraints of ’=’ horizontal, ’=’ vertical, ’x’ horizontal and ’x’ vertical. For the ’=’ terms, we need the energy when they are equal to be zero. We achieve this by knowing that, for a binary case $x(1-y)+(1-x)y$ is null if $x=y$ and 1 if $x\neq y$. We need the second summand so that there is no null if $x=0$ and $y=1$. Thus, the terms for equals are
$Q_{He}=\sum_{k=1}^{N_{He}}$ $\left(x_{I_{k}^{He},J_{k}^{He}}\left(1-x_{I_{k}^{He},J_{k}^{He}+1}\right)+\right.$
$+$ $\left.\left(1-x_{I_{k}^{He},J_{k}^{He}}\right)x_{I_{k}^{He},J_{k}^{He}+1}\right),$ (42)
$Q_{Ve}=\sum_{k=1}^{N_{Ve}}$ $\left(x_{I_{k}^{Ve},J_{k}^{Ve}}\left(1-x_{I_{k}^{Ve}+1,J_{k}^{Ve}}\right)+\right.$
$+$ $\left.\left(1-x_{I_{k}^{Ve},J_{k}^{Ve}}\right)x_{I_{k}^{Ve}+1,J_{k}^{Ve}}\right).$ (43)
For the case ’x’ we need a term that is null only when both values are different. To do so, we will take advantage of the fact that $xy+(1-x)(1-y)$ is equal to 0 if $x\neq y$ and 1 if $x=y$. Thus, the terms are
$Q_{Hc}=\sum_{k=1}^{N_{Hc}}$ $\left(x_{I_{k}^{Hc},J_{k}^{Hc}}x_{I_{k}^{Hc},J_{k}^{Hc}+1}+\right.$
$+$ $\left.\left(1-x_{I_{k}^{Hc},J_{k}^{Hc}}\right)\left(1-x_{I_{k}^{Hc},J_{k}^{Hc}+1}\right)\right),$ (44)
$Q_{Vc}=\sum_{k=1}^{N_{Vc}}$ $\left(x_{I_{k}^{Vc},J_{k}^{Vc}}x_{I_{k}^{Vc}+1,J_{k}^{Vc}}+\right.$
$+$ $\left.\left(1-x_{I_{k}^{Vc},J_{k}^{Vc}}\right)\left(1-x_{I_{k}^{Vc}+1,J_{k}^{Vc}}\right)\right).$ (45)
Given this formulation, we can perform the following optimization of variables by taking advantage of the symbols ’=’ and ’x’. As we know, if there is one of these symbols between two cells, both variables must have the same or different value. In this case in which we can only have two possible values, this implies that these two variables are going to be completely dependent, so that if we determine one, the other is automatically determined. Therefore, these last four terms of the formulation are unnecessary if we reduce the number of variables as follows.
In our optimization, we optimize only one of the variables of the pair affected by the symbols, and in case we have a chain of variables all affected by each other, we optimize only one of them. The rule we follow is that in each horizontal pair, we optimize the variable on the left, and in each vertical pair, the one on top. In case of having whole regions affected, we choose the variable that is more to the left, and in case of having several in the same column, the one that is higher among them.
To take this dependence into account, we will make the following substitution in the first four terms:
  • In the case of having an ’=’ between $(i,j)$ and $(i,j+1)$
    $x_{i,j+1}=x_{i,j}.$ (46)
  • In the case of having an ’=’ between $(i,j)$ and $(i+1,j)$
    $x_{i+1,j}=x_{i,j}.$ (47)
  • In the case of having an ’x’ between $(i,j)$ and $(i,j+1)$
    $x_{i,j+1}=(1-x_{i,j}).$ (48)
  • In the case of having an ’x’ between $(i,j)$ and $(i+1,j)$
    $x_{i+1,j}=(1-x_{i,j}).$ (49)
In this way, we are able to impose this constraint automatically, without having to optimize it, reducing the number of variables to optimize. Thus, the number of variables to be optimized will be, at the most
$NM-N_{I0}-N_{I1}-N_{He}-N_{Hc}-N_{Ve}-N_{Vc}.$
Finally, the global non-repetition condition cannot be written as a QUBO term, so we will have to be confident that the solutions to the problem do not conflict with this constraint.
Therefore, the QUBO formulation of the TTP is
$Q=$ $Q_{HR}+Q_{VR}+Q_{Hr}+Q_{Vr}=$
$=$ $\sum_{i,j=1}^{N,M-2}\left(\frac{3}{2}-\sum_{k=j}^{j+2}x_{ik}\right)^{2}+$
$+$ $\sum_{i,j=1}^{N-2,M}\left(\frac{3}{2}-\sum_{k=i}^{i+2}x_{kj}\right)^{2}+$
$+$ $\sum_{i=1}^{N}\left(\frac{M}{2}-\sum_{j=1}^{M}x_{ij}\right)^{2}+$
$+$ $\sum_{j=1}^{M}\left(\frac{N}{2}-\sum_{i=1}^{N}x_{ij}\right)^{2}.$ (50)
Thus, a solution to the problem that satisfies the constraints will have an energy of $\frac{1}{2}(NM-N-M)$.
3.3 Generalizations of the TTP
Having solved the original problem, we will present several generalizations of TTP.
3.3.1 Unbalanced regularity condition
The first generalization we propose is to allow for each row or column to have a different number of zeros than ones. Since we know that the number of ones determines the number of zeros, we can say that if there are $N$ cells in that direction, we want there to be $N_{1}$ ones and $N-N_{1}$ zeros. This can also change between rows and columns. We say that in row $i$ we want $M_{1,i}$ ones, and in column $j$ we want $N_{1,j}$ ones. These generalized terms, for both horizontal and vertical, are
$Q_{Hr}=\sum_{i=1}^{N}\left(M_{1,i}-\sum_{j=1}^{N}x_{ij}\right)^{2},$ (51)
$Q_{Vr}=\sum_{j=1}^{M}\left(N_{1,j}-\sum_{i=1}^{N}x_{ij}\right)^{2}.$ (52)
3.3.2 Diagonal repetition
The second generalization we propose is that, in addition to not being able to have more than two ones or zeros in a row vertically or horizontally, it should not be possible to have more than two ones or zeros diagonally. For this, we define 2 new terms, exactly the same as the original ones for this, but changing the second summation, so that they evaluate the lower diagonals for each cell. We have $Q_{DLR}$ for the left diagonals and $Q_{DRR}$ for the right diagonals:
$Q_{DLR}=\sum_{i,j=1,3}^{N-2,M}\left(\frac{3}{2}-\sum_{k=0}^{2}x_{i+k,j-k}\right)^{2},$ (53)
$Q_{DRR}=\sum_{i,j=1}^{N-2,M-2}\left(\frac{3}{2}-\sum_{k=0}^{2}x_{i+k,j+k}\right)^{2},$ (54)
being $x_{a,b}=0\quad\forall b<1,a\in[1,N]$, and $\forall b>M,a\in[0,N-1]$.
3.3.3 Regions condition
Another possible generalization is the imposition of a specific number of ones in certain colored regions. That is, for a set of $N_{R}$ regions, in the $k$-th region $R_{k}$ there must be $m_{k}$ ones, as we can see in Fig. 8 a. This condition is imposed with a term
$\sum_{k=1}^{N_{R}}\left(m_{k}-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2}.$ (55)
Figure 8: Generalized cases. a) Takuzu with colored regions, b) toroidal Takuzu, where the blue cells have fixed initial values and the red cells are a consequence of the repetition constraint.
3.3.4 Toroidal board
An interesting generalization is to consider a toroidal board, so that the condition that there are no more than 2 ones or zeros in a row vertically or horizontally is cyclically extended to the other end of the board, as can be seen in Fig. 8 b.
The only term that will change will be the repetition constraint, so that we will have to perform the summations with all rows and columns, making $x_{a,b}=x_{a-N,b}$ if $a>N$ and $x_{a,b}=x_{a,b-M}$ if $b>M$. In this way we will take into account the toroid shape. In case we consider a cylinder instead of a toroid, we will only have to keep one of these changes, in the direction that is cyclic.
3.3.5 Long distance equal/cross
In Tango’s original problem, we have that there can only be ’=’ and ’x’ symbols acting between vertically and horizontally adjacent cells. We extend this idea to mean that both symbols can act between any two cells, as can be seen in Fig. 9 a. To implement this generalization, we need only allow that the variable changes exhibited in (46), (47), (48) and (49) can be made between two non-adjacent variables. The rest of the formulation is the same.
On the other hand, we also generalize the possibility of defining regions in which all their variables have to be equal (equality regions), and separately defining that a pair of equality regions have to have different values for their variables from each other, as in Fig. 9 b. To do this, just impose a set of ’=’ symbols between the cells of the region, so that they form a chain of equalities with each other, and then impose a single ’x’ symbol between any pair of cells of one and the other region. In this way, since all cells in each region have to be equal, if a cell in one is different from a cell in the other, all cells in one region will be opposite cells in the other region.
Figure 9: a) Symbols that act over long distances. b) Regions of equality that are different from each other. All cells in the blue region have the same value, all cells in the yellow region have the same value, and the blue cells all have different values than the yellow cells.